Lazorwolf64:

Help y=5x-7 -3x-2y=-12

Joe348:

Substitute 5x-7 for y in the equation -3x-2y=-12 \[-3x-2(5x-7)=-12\] \[-13x+14=-12\] Add 14 to both sides \[-13x=-26\] Divide both sides by -13 \[x=2\]

Joe348:

Substitute 2 for x in the equation y=5x-7 \[(5)(2)-7\] \[y=3\]

Joe348:

\[x=2 \] \[y=3\]

Lazorwolf64:

Thank you!!!!!

Joe348:

glad to help<333